Solve the equation by graphing the function
2x^2=3x^2-2x-8
show all work please thanks!

Answer by ewatrrr(24785)

y = x^2 - 2x - 8 -8 the y-intercept
y = (x-1)^2 - 1 - 8
y = (x-1)^2 - 9. Parabola Opening Upward: V(1,-9) x = 1 the Line of Symmetry